# Dynamic planning in hierarchical active inference

### v2

@@ 2.1 A simple agent
Figure 2: (a) In this task, the agent (a single DoF) has to reach a target angle represented by the red circle. Estimated and real arms are displayed in cyan and blue, respectively. Here, $\bm{\pi}_{\eta}=0$, $\bm{\rho}=120$°, and $\bm{\mu}$ was initialized to $-40$°. The time step is indicated in the bottom left corner of each frame. Since the belief was initialized at a negative value, the likelihood initially pulls the arm toward the wrong direction before adapting to the dynamics attractor. (b) The top graph shows the evolution of the real angle $\bm{x}$, the belief over it $\bm{\mu}$, and the target angle $\bm{\rho}$. The middle graph shows the evolution of the belief over the velocity $\bm{\mu}^{\prime}$ and the belief derivative $\dot{\bm{\mu}}$. The bottom graph shows the evolution of all the components that comprise the belief update: the belief over the velocity $\bm{\mu}^{\prime}$, the likelihood gradient $\partial_{\mu}\bm{g}^{T}\bm{\pi}_{o}\bm{\varepsilon}_{o}$, the dynamics gradient $\partial_{\mu}\bm{f}^{T}\bm{\pi}_{x}\bm{\varepsilon}_{x}$, and the weighted dynamics prediction error $-\bm{\pi}_{x}\bm{\varepsilon}_{x}$. The latter has been plotted for better comparing its magnitude with the other components, although affecting the 1st temporal order.
@@
The interactions between these quantities can be better understood from Figure 2, showing a reaching movement with the defined dynamics function and the trajectories of the agent’s generative model. Here, the belief is subject to two different forces: a likelihood gradient pushing it toward what it is currently perceiving (i.e., the real angle), and the other components that steer it toward the biased dynamics (i.e., the target angle $\bm{\rho}$). Note how in the third plot, of the three components that comprise the belief update, the backward error $\partial_{\mu}\bm{f}^{T}\bm{\pi}_{x}\bm{\varepsilon}_{x}$ has the least impact in the overall direction of update. While the exact interactions arising from the dynamics prediction error have yet to be analyzed, in the following we assume that goal-directed behavior is achieved through the forward error at the 1st order $-\bm{\pi}_{x}\bm{\varepsilon}_{x}$. An alternative would be to directly control the backward error without maintaining a belief over increasing temporal orders [85], which however requires to take a gradient into account and may be more challenging when defining appropriate attractors to reach a goal. Finally see how, in the middle plot, the agent tries at every instant to minimize the difference between $\bm{\mu}^{\prime}$ and $\dot{\bm{\mu}}$, thus tracking the actual path of the hidden states.
@@
But how does this agent actually move? As mentioned in the Introduction, action is the other side of the coin of the free energy principle, through which the agent samples those observations that conform to its prior beliefs. In fact, in addition to the perceptual inference typical of predictive coding, active inference assumes that organisms minimize free energy also by interacting with the environment; this minimization breaks down to an even simpler update that only depends on observation prediction errors $\bm{\varepsilon}_{o}$. Since these prediction errors are generated from the agent’s belief, this means that whenever the latter is biased toward some preferred state, movement naturally follows. There is thus a delicate balance between perception – in which prediction errors climb up the hierarchy to bring the belief closer to the observations – and action – in which prediction errors are suppressed at a low level so that the observations are brought closer to their predictions. However, there is an open issue regarding how active inference should be practically realized in continuous time. A few studies demonstrated that using exteroceptive information directly for computing motor commands could result in smoother movements and resolution of visuo-proprioceptive conflicts [28, 41, 43], and in fact some robotic implementations effectively used this approach [85, 86]. However, evidence seems to indicate that motor commands are generated by suppression of proprioceptive information only [34, 33], which is already in the intrinsic reference frame needed for movement and thus results in an easier inverse dynamics. For this reason, in the following we assume that – indicating by the subscript $p$ the proprioceptive domain – movements are realized by minimizing the free energy with respect to proprioceptive prediction errors $\bm{\varepsilon}_{p}$:
$\dot{\bm{a}}=-\partial_{a}\mathcal{F}=-\partial_{a}\bm{g}_{p}^{T}\bm{\pi}_{p}\bm{\varepsilon}_{p}$ (9)
@@
where $\partial_{a}\bm{g}_{p}$ performs an inverse dynamics from the proprioceptive predictions to the motor commands $\bm{a}$, likely to be implemented by classical spinal reflex arcs. As a last note, the actions can also depend on multiple orders – velocity, acceleration, and so on – allowing more efficient movement and control [89, 90, 91, 92], but since it is beyond our scope, we only minimize the 0th order. Nonetheless, 1st-order movements – e.g., maintaining a constant velocity – are still possible by specification of appropriate dynamics of the hidden states.
@@ 2.2 Tracking objects
Figure 3: The target is now encoded in the hidden causes $\bm{v}$, generating a dynamic attractor for object tracking. In fact, both hidden states and hidden causes generate predictions through parallel likelihood functions $\bm{g}_{x}$ and $\bm{g}_{v}$, and both concur in estimating the 1st-order hidden states $\bm{x}^{\prime}$.
The simple agent defined in the previous section can only realize fixed trajectories embedded in the dynamics function, so how can we let it track moving objects? This is usually done by introducing a key concept in active inference, the hidden causes $\bm{v}$, which link hierarchical levels and specify how the dynamics function evolves. In the active inference literature of motor control, they are also used to encode the target to be reached [28, 67, 74, 93], as depicted in Figure 3. Considering the target as a causal variable for the hidden states and sensory observations makes sense from an active perspective whereby “it is an object I want to reach that generates my movements”. Now, the agent’s generative model becomes:
$p(\tilde{\bm{x}},\bm{v},\bm{o})=p(\bm{o}|\bm{x},\bm{v})p(\bm{x})p(\bm{x}^{\prime}|\bm{x},\bm{v})p(\bm{v})$ (10)
where:
[…]
Figure 4: (a) In this task, the agent has to track a target angle rotating at a constant velocity. Estimated and real targets are displayed in purple and red, respectively. Here, $\bm{\pi}_{\eta,x}=0$, $\bm{\pi}_{\eta,v}=0$, $\bm{v}$ was initialized to $60$°, and both $\bm{\mu}$ and $\bm{\nu}$ were initialized to 00°. Here, the belief over the hidden causes pulls the belief over the hidden states with it while approaching the real target angle. (b) The top graph shows the evolution of the real angle $\bm{x}$, the belief over it $\bm{\mu}$, the target angle $\bm{v}$, and the belief over it $\bm{\nu}$. The middle graph shows the evolution of the belief over the velocity $\bm{\mu}^{\prime}$ and the belief derivative $\dot{\bm{\mu}}$, as before. The bottom graph shows the evolution of all the components that comprise the hidden causes update: the likelihood gradient $\partial_{\nu}\bm{g}_{v}^{T}\bm{\pi}_{o,v}\bm{\varepsilon}_{o,v}$, and the dynamics gradient $\partial_{\nu}\bm{f}^{T}\bm{\pi}_{x}\bm{\varepsilon}_{x}$. Note how, in the middle plot, the estimated 1st temporal order stabilizes to a non-zero value as the agent rotates with a constant angular velocity.
Note that there are two priors, one over the hidden states and another over the hidden causes, respectively denoted by $\bm{\eta}_{x}$ and $\bm{\eta}_{v}$. Note also that both dynamics and likelihood functions depend on the hidden causes, and that we assumed a further factorization for the likelihood:
[…]
where $\bm{o}_{x}$ and $\bm{o}_{v}$ denote the hand and target observations, respectively. It is this additional connection between hidden causes and observations that makes the agent able to operate in dynamic environments. In fact, we can define the following dynamics function:
@@
where we just replaced the static target $\bm{\rho}$ with the hidden causes. Then, the posterior belief over the hidden causes $\bm{\nu}$ is updated according to:
$\dot{\bm{\nu}}=-\partial_{\nu}\mathcal{F}=-\bm{\pi}_{\eta_{v}}\bm{\varepsilon}_{\eta_{v}}+\partial_{\nu}\bm{g}_{v}^{T}\bm{\pi}_{o,v}\bm{\varepsilon}_{o,v}+\partial_{\nu}\bm{f}^{T}\bm{\pi}_{x}\bm{\varepsilon}_{x}$ (14)
@@
where we defined the following observation and prior prediction errors:
$\begin{split}\bm{\varepsilon}_{o,v}&=\bm{o}_{v}-\bm{g}_{v}(\bm{\nu})\\ \bm{\varepsilon}_{\eta_{v}}&=\bm{\nu}-\bm{\eta}_{v}\end{split}$ (15)
@@
As evident, the hidden causes are subject to a prior prediction error, a backward dynamics error, and a backward likelihood error – similar to the update of the hidden states, with the only difference that this kind of inference is over a state and not a path. Via the backward likelihood error, the agent can correctly estimate the target configuration whenever it moves, as shown in the tracking simulation of Figure 4. Concerning the dynamics prediction error, it can now flow into two different pathways: specifically, the role of the gradients $\partial_{\mu}\bm{f}$ and $\partial_{\nu}\bm{f}$ are to respectively infer the positional state and the cause that may have generated a particular velocity; their actual role will be clear in Chapter 4.
2.3 Intention modulation and multi-step behavior
@@
Although capable of operating in dynamic contexts, the last approach still portrays a simple scenario in which a specific target has no internal dynamics and has always the role of a cause for a hidden state. In other words, it does not permit modeling realistic tasks such as a pick-and-place operation, where an object can either be the cause of a reaching and grasping movement, or the consequence of another cause such as a goal position, resulting in a placing movement; critically, it does not allow to model a task wherein not only the dynamics of the self, but also the dynamics of the target must be learned (e.g., if a moving object should be grasped on the fly, the agent should infer its trajectory to anticipate where it will fall).
@@
It follows that to operate in a complex environment, the agent must (i) maintain complete representations for each entity that it wants to interact with, and (ii) flexibly assign causes and consequences for the next movement depending on the current context – in a similar way to policies in discrete models, as will be explained later. Therefore, we first encode multiple environmental entities in the hidden states, i.e., $\bm{x}=[\bm{x}_{1},\dots,\bm{x}_{N}]$, where $N$ is the number of entities [43]. Consequently, the factorized likelihood function generates specular observations for each element:
$\bm{o}=[\bm{o}_{1},\dots,\bm{o}_{N}]=[\bm{g}_{1}(\bm{x}_{1}),\dots,\bm{g}_{N}(\bm{x}_{N})]$ (16)
@@
This structure is similar to the previous model, except that the target is now embedded in the hidden states along with the hand, and that there is no connection between hidden causes and observations. We could define a similar factorization for the hidden causes and dynamics function, i.e., $\bm{x}^{\prime}=[\bm{f}_{1}(\bm{x}_{1},\bm{v}_{1}),\dots,\bm{f}_{N}(\bm{x}_{N},\bm{v}_{N})]$, such that each entity would have an independent dynamics biased by a specific cause (e.g., where the hand or the target will be in the future); however, this is of limited use in a pick-and-place operation that demands interaction between entities. We hence compute a potential hidden state with a single function, such as:
$\bm{i}(\bm{x})=\bm{W}\bm{x}+\bm{b}$ (17)
@@
The weights $\bm{W}$ perform a linear transformation of the hidden states that combines every entity, while the bias $\bm{b}$ imposes a static configuration over them [44]. Equation 17 can be realized through simple neural connections, wherein the weights are encoded as synaptic strengths and the bias represents the threshold needed to fire a spike. An error is then computed between this potential state and the current one:
$\bm{e}_{i}=\bm{i}(\bm{x})-\bm{x}$ (18)
@@
multiplying the error by a single-value hidden cause $v$. Thus, the latter is not intended as an explicit trajectory prior over the hidden states (e.g., encoding where my hand will be in the future), whose role is now delegated to the bias $\bm{b}$; but as an attractor gain, whereby a high value implies a strong force toward the potential state. As a result, we have an additional modulation that combines with the dynamics precision $\bm{\pi}_{x}$; their interactions will be explained in Chapter 4. Since $\bm{i}(\bm{x})$ is used to define a path for the current hidden states aiming to produce a desired configuration, we call it an intention. Similarly, we refer to $\bm{e}_{i}$ as an intention prediction error; note however that this quantity is not strictly a prediction error, although it would be possible to design the model to call it as such.
@@
Figure 5: (a) Factor graph of the unit with environmental beliefs and dynamic behavior. Each variable is factorized into independent components that encode the agent’s bodily states and external physical variables, e.g., objects to interact with. A single dynamics function combines the environmental beliefs to compute a specific intention. The hidden causes now define an attractor gain expressing the strength of the intention. The weights $\bm{W}$ of the intention can be used, e.g., to track moving objects, while the bias $\bm{b}$ realizes a static configuration. See [43, 75] for more details. (b) Factor graph of the unit with multiple intentions. Every hidden cause $v_{m}$, coupled with an independent intention of the dynamics function, is combined to produce a dynamic trajectory $\bm{\eta}_{x}^{\prime}$, comprising all the environmental beliefs $\bm{x}_{n}$. The transition between intentions can be achieved by a higher-level prior, e.g., a belief over tactile sensations.
@@
In summary, as shown in Figure 5(a), the dynamics function is not composed of segregated pathways as for the likelihood, but affects all the environmental entities at once – e.g., it computes a trajectory for the hand depending on the target. The steps performed by the agent during a reaching movement are the following: (i) the 0th order imposes a dynamic trajectory to the 1st order and generates a sensory prediction; (ii) the 0th order infers the consequences of its predictions, hence it is now biased toward both the intention and the observation; (iii) a proprioceptive prediction is generated from this new biased position, eventually driving action. This approach can be seen as a generalization of [74] where, in the context of oculomotor behavior, the target and the center of gaze were encoded as hidden states, with their own dynamics and attracted by a hidden location. Although somewhat limited compared to non-linear dynamics functions (e.g., obstacle avoidance can be realized by dynamics functions built from repulsive potentials [44]), through the specific form defined above – and along with the hidden states factorization – there is a high flexibility in designing intentions for complex interactions. Further, interpreting the hidden causes as a gain still makes sense from an active inference perspective because what is represented at a higher level is the intention to move at the target, while the target location is inferred at a lower level.
@@
Taken alone, considering a hidden cause as an attractor gain may not seem so helpful. However, as depicted in Figure 5(b), we can combine $M$ intentions in the following way:
$\begin{split}\bm{\eta}_{x,m}^{\prime}&=\bm{f}_{m}(\bm{x},\bm{v})=v_{m}\bm{e}_{i,m}\\ \bm{\eta}_{x}^{\prime}&=\bm{f}(\bm{x},\bm{v})=\sum_{m}^{M}\bm{\eta}_{x,m}^{\prime}\end{split}$ (20)
@@
In short, trajectories $\bm{\eta}_{x,m}^{\prime}$ are separately computed from each intention $\bm{i}_{m}$ and with their respective gains; then, the final trajectory $\bm{\eta}_{x}^{\prime}$ is found by combining all of them. The reason why we used the prior notation for the trajectory predictions will be clear in Chapter 4. Note for the moment that, as before, there is a different structure compared to the likelihood. While an observation is generated through a parallel pathway for every environmental belief, each function $\bm{f}_{m}(\bm{x},\bm{v})$ combines all of them in a specific way. Since each intention prediction error is proportional to its hidden cause $v_{v}$, the latter lends itself to a parallelism with the policies of discrete models, as will be explained later: if $v_{m}$ is set to $1$ and all the others to 00, the hidden states will be subject only to intention $m$; conversely, if multiple hidden causes are active, the hidden states will be pulled toward a combination of the corresponding intentions. This means that the hidden causes act both as attractor gains – expressing the absolute strength by which the belief is steered toward the desired direction – and as intention modulators – defining the relative strength between each desired state. The resulting dynamics prediction error:
$\bm{\varepsilon}_{x}=\bm{\mu}^{\prime}-\bm{\eta}_{x}^{\prime}$ (21)
will then realize an average trajectory that the agent predicts for the current situation. This approach is effective for two reasons. First, it allows defining a composite movement in terms of simpler subgoals, which can be tackled separately; this can be helpful, for instance, if one has to analyze the behavior of an agent when subject to two or more opposing priors [43]. But the main utility is that a fixed multi-step behavior can be achieved [75] which does not need to modify the dynamics function at each step but only to modulate the hidden causes, since the model already encodes all the intermediate goals that the agent will pass through. Transitions between continuous trajectories could then be realized by higher-level priors over the hidden causes, e.g., a belief over tactile sensations for multi-step reaching, such as in the simulation of Figure 6.

### v3

@@ 2.1 A simple agent
Figure 2: (a) In this task, the agent (a single DoF) has to reach a target angle represented by the red circle. Estimated and real arms are displayed in cyan and blue, respectively. Here, $\bm{\pi}_{\eta,x}=0$, $\bm{\rho}=120$°, and $\bm{\mu}_{x}$ was initialized to $-40$°. The time step is indicated in the bottom left corner of each frame. Since the belief was initialized at a negative value, the likelihood initially pulls the arm toward the wrong direction before adapting to the dynamics attractor. (b) The top graph shows the evolution of the real angle $\bm{x}$, its belief $\bm{\mu}_{x}$, and the target angle $\bm{\rho}$. The middle graph shows the evolution of the belief of the velocity $\bm{\mu}_{x}^{\prime}$ and the belief derivative $\dot{\bm{\mu}}_{x}$. The bottom graph shows the evolution of all the components that comprise the belief update: the belief of the velocity $\bm{\mu}_{x}^{\prime}$, the likelihood gradient $\partial_{x}\bm{g}^{T}\bm{\Pi}_{o,p}\bm{\varepsilon}_{o,p}$, the dynamics gradient $\partial_{x}\bm{f}^{T}\bm{\Pi}_{x}\bm{\varepsilon}_{x}$, and the weighted dynamics prediction error $-\bm{\Pi}_{x}\bm{\varepsilon}_{x}$. The latter has been plotted to compare its magnitude with the other components, although affecting the 1st temporal order.
@@
The interactions between these quantities can be better understood from Figure 2, showing a reaching movement with the defined dynamics function and the trajectories of the agent’s generative model. Here, the belief is subject to two different forces: a likelihood gradient pushing it toward what it is currently perceiving (i.e., the real angle), and the other components that steer it toward the biased dynamics (i.e., the target angle $\bm{\rho}$). Note how in the third plot, of the three components that comprise the belief update, the backward error $\partial_{x}\bm{f}^{T}\bm{\Pi}_{x}\bm{\varepsilon}_{x}$ has the smallest amplitude. While the exact interactions arising from the dynamics prediction error have yet to be analyzed, in the following we assume that goal-directed behavior is achieved through the forward error at the 1st order $-\bm{\Pi}_{x}\bm{\varepsilon}_{x}$. An alternative would be to directly control the backward error without maintaining a belief of the 1st order [70], which however requires taking a gradient into account and may be more challenging when defining appropriate attractors. Finally see how, in the middle plot, the agent tries at every instant to minimize the difference between $\bm{\mu}_{x}^{\prime}$ and $\dot{\bm{\mu}}_{x}$, thus tracking the actual path of the hidden states.
@@
But how does the agent move in practice? As mentioned in the Introduction, action is the other side of the coin of the free energy principle, through which the agent samples those observations conforming to its prior beliefs. In fact, in addition to the perceptual inference typical of predictive coding, active inference assumes that organisms minimize free energy also by interacting with the environment; this minimization breaks down to an even simpler update that only depends on (proprioceptive) prediction errors $\bm{\varepsilon}_{o,p}$. Since these prediction errors are generated from the agent’s belief, this means that whenever the latter is biased toward some preferred state, movement naturally follows. There is thus a delicate balance between perception – in which prediction errors climb up the hierarchy to bring the belief closer to the observations – and action – in which prediction errors are suppressed at a low level so that the observations are brought closer to their predictions. However, there is an open issue regarding how active inference should be practically realized in continuous time. A few studies demonstrated that using exteroceptive information directly for computing motor commands could result in smoother movements and resolution of visuo-proprioceptive conflicts [32, 44, 64], and in fact some robotic implementations effectively used this approach [70, 71]. However, evidence seems to indicate that motor commands are generated by suppression of proprioceptive signals only [38, 37], which is already in the intrinsic reference frame needed for movement and thus results in easier inverse dynamics. For this reason, in the following we assume that movements are realized by minimizing the free energy with respect to proprioceptive prediction errors:
$\dot{\bm{a}}=-\partial_{a}\mathcal{F}=-\partial_{a}\bm{g}_{p}^{T}\bm{\Pi}_{o,p}\bm{\varepsilon}_{o,p}$ (7)
@@
where $\partial_{a}\bm{g}_{p}$ performs an inverse dynamics from proprioceptive predictions to motor commands $\bm{a}$, likely to be implemented by classical spinal reflex arcs. As a last note, the actions can also depend on multiple orders – velocity, acceleration, and so on – allowing more efficient movement and control [74, 75, 76, 77], but since it is beyond our scope, we only minimize the 0th order. Nonetheless, 1st-order movements – e.g., maintaining a constant velocity – are still possible by specifying appropriate dynamics of the hidden states.
@@ 2.2 Tracking objects
Figure 3: (a) The target is now encoded in the hidden causes $\bm{v}$, generating a dynamic attractor for object tracking. In fact, both hidden states and hidden causes generate predictions through proprioceptive and visual likelihood functions $\bm{g}_{p}$ and $\bm{g}_{v}$, and both concur in estimating the 1st-order hidden states $\bm{x}^{\prime}$. (b) Agent’s generative model.
The simple agent defined in the previous section can only realize fixed trajectories embedded in the dynamics function, so how can it track moving objects? This is usually done by introducing a key concept in active inference, the hidden causes $\bm{v}$, which link hierarchical levels and specify how the dynamics function evolves. In the active inference literature of motor control, they are also used to encode the target to be reached [32, 78, 63, 79], as depicted in Figure 3(a). Considering the target as a causal variable for the hidden states and sensory observations makes sense from an active perspective whereby “it is an object I want to reach that generates my movements”. Now, the agent’s generative model becomes that of Figure 3(b). Note that there are two priors, one over the hidden states and another over the hidden causes, respectively denoted by $\bm{\eta}_{x}$ and $\bm{\eta}_{v}$. Also, both dynamics and likelihood functions depend on the hidden causes, and we assumed a further factorization for the likelihood, where $\bm{o}_{p}$ and $\bm{o}_{v}$ denote the arm and target observations, respectively. For simplicity, we assume that the visual likelihood function $\bm{g}_{v}$ is a simple identity generating the target angle. It is through the connection between hidden causes and observations that the agent can operate in dynamic environments. In fact, we can define the following dynamics function:
@@
Figure 4: (a) In this task, the agent has to track a target angle rotating at a constant velocity. Estimated and real targets are displayed in purple and red, respectively. Here, $\bm{\Pi}_{\eta,x}=0$, $\bm{\Pi}_{\eta,v}=0$, $\bm{v}$ was initialized to $60$°, and both $\bm{\mu}_{x}$ and $\bm{\mu}_{v}$ were initialized to 00°. Here, the belief of the hidden causes pulls the belief of the hidden states with it while approaching the real target angle. (b) The top graph shows the evolution of the real angle $\bm{x}$, its belief $\bm{\mu}_{x}$, the target angle $\bm{v}$, and its belief $\bm{\mu}_{v}$. The middle graph shows the evolution of the belief of the velocity $\bm{\mu}_{x}^{\prime}$ and the belief derivative $\dot{\bm{\mu}}_{x}$, as before. The bottom graph shows the evolution of all the components that comprise the hidden causes update: the likelihood gradient $\partial_{v}\bm{g}_{v}^{T}\bm{\Pi}_{o,v}\bm{\varepsilon}_{o,v}$, and the dynamics gradient $\partial_{v}\bm{f}^{T}\bm{\Pi}_{x}\bm{\varepsilon}_{x}$. Note how, in the middle plot, the estimated 1st temporal order stabilizes to a non-zero value as the agent rotates with a constant angular velocity.
@@
where we just replaced the static target $\bm{\rho}$ with the hidden causes. As for the hidden states, we define a (Gaussian) approximate posterior over the hidden causes $\bm{q}(\bm{v})=\mathcal{N}(\bm{\mu}_{v},\bm{P}_{v})$, with mean $\bm{\mu}_{v}$ and precision $\bm{P}_{v}$. Then, the mean is updated according to:
$\dot{\bm{\mu}}_{v}=-\partial_{v}\mathcal{F}=-\bm{\Pi}_{\eta,v}\bm{\varepsilon}_{\eta,v}+\partial_{v}\bm{g}_{v}^{T}\bm{\Pi}_{o,v}\bm{\varepsilon}_{o,v}+\partial_{v}\bm{f}^{T}\bm{\Pi}_{x}\bm{\varepsilon}_{x}$ (9)
@@
where we defined the following observation and prior prediction errors in terms of the approximate posterior:
$\begin{split}\bm{\varepsilon}_{o,v}&=\bm{o}_{v}-\bm{g}_{v}(\bm{\mu}_{v})\\ \bm{\varepsilon}_{\eta,v}&=\bm{\mu}_{v}-\bm{\eta}_{v}\end{split}$ (10)
@@
As evident, the hidden causes are subject to a prior prediction error, a backward dynamics error, and a backward likelihood error – similar to the update of the hidden states, except that this inference is over a state and not a path. Via the backward likelihood error, the agent can correctly estimate the target configuration whenever it moves, as shown in the tracking simulation of Figure 4. Concerning the dynamics prediction error, it can now flow into two different pathways: specifically, the role of the gradients $\partial_{x}\bm{f}$ and $\partial_{v}\bm{f}$ are to respectively infer the state and the cause that may have generated a particular velocity; their actual role will be clear in Chapter 4.
2.3 Intention modulation and object affordances
@@
Although capable of operating in dynamic contexts, the last approach still reproduces a simple scenario in which a target has no internal dynamics and always has the role of a cause for a hidden state. In other words, it does not permit modeling realistic tasks such as a pick-and-place operation, where an object is first the cause of a reaching and grasping movement, but then it becomes the consequence of another cause such as a goal position, resulting in a placing movement; critically, it does not allow to model a task wherein not only the dynamics of the self, but also the dynamics of the target must be learned (e.g., if a moving object should be grasped on the fly, the agent should infer its trajectory to anticipate where it will fall).
@@
It follows that to operate in a complex environment, the agent must (i) maintain complete representations for each entity that it wants to interact with, and (ii) flexibly assign causes and consequences for the next movement depending on the current context – in a similar way to policies in discrete models, as will be explained later. Therefore, we first encode multiple environmental entities as potential body configurations in the hidden states, i.e., $\bm{x}=[\bm{x}_{0},\bm{x}_{1},\dots,\bm{x}_{N}]$, where $\bm{x}_{0}$ is the actual body configuration (as before), and $N$ is the number of entities [64]. Consequently, the factorized likelihood function generates a proprioceptive observation for the first component $\bm{x}_{0}$, and visual observations for each entity:
$\bm{o}=[\bm{o}_{p},\bm{o}_{v,1},\dots,\bm{o}_{v,N}]=[\bm{g}_{p}(\bm{x}_{0}),\bm{g}_{v}(\bm{x}_{1}),\dots,\bm{g}_{v}(\bm{x}_{N})]$ (11)
@@
Here, visual observations are assumed to be in the Cartesian domain, so the visual likelihood function $\bm{g}_{v}$ generates the hand positions of the potential configurations through forward kinematics. This structure is similar to the previous model, except that the target is now embedded in the hidden states along with the hand, and that there is no connection between hidden causes and observations. We could define a similar factorization for the hidden causes and dynamics function, such that each entity would have an independent dynamics biased by a specific cause (e.g., where the arm or the target will be in the future); however, this is of limited use in a pick-and-place operation that demands interaction between entities. We hence compute an intentional state with a single function, such as:
$\bm{i}_{m}(\bm{x})=\bm{W}_{m}\bm{x}+\bm{b}_{m}$ (12)
@@
The weights $\bm{W}_{m}$ perform a linear transformation of the hidden states that combines every entity, while the bias $\bm{b}_{m}$ imposes a static configuration [80]. Equation 12 can be realized through simple neural connections, wherein the weights are encoded as synaptic strengths and the bias represents the threshold needed to fire a spike, and both can be known a priori or learned via sensory evidence. An error is then computed between this intentional state and the current one:
$\bm{e}_{i,m}=\bm{i}_{m}(\bm{x})-\bm{x}$ (13)
@@
multiplying the error by a single-value hidden cause $v_{m}$. Thus, the latter is not intended as an explicit trajectory prior over the hidden states (e.g., encoding where my arm will be in the future), whose role is now delegated to the bias $\bm{b}_{m}$; but as an attractor gain, whereby a high value implies a strong force toward the future state. Since $\bm{i}_{m}$ is used to define a path for the hidden states aiming to produce a desired configuration, we call it an intention.
@@
Figure 5: (a) Factor graph of the unit with object affordances. Hidden states are factorized into independent components that encode the actual bodily states and potential configurations related to the objects. The first component $\bm{x}_{0}$ generates proprioceptive predictions, while the successive components generate visual predictions of the objects. Every hidden cause $v_{m}$ now defines an attractor gain expressing the strength of an agent’s intention (encoded as a distinct evolution of the world). These are combined to produce a trajectory $\bm{\eta}_{x}^{\prime}$, comprising all the body configurations $\bm{x}_{n}$. The transition between intentions can be achieved by a higher-level prior, e.g., a belief of tactile sensations. The weights $\bm{W}_{m}$ of the intention can be used, e.g., to track moving objects, while the bias $\bm{b}_{m}$ realizes a static configuration. See [64, 66] for more details. (b) Agent’s generative model.
@@
The dynamics function of Equation 14 is not composed of segregated pathways as for the likelihood, but affects all the environmental entities at once – e.g., it computes a trajectory for the arm depending on the target. The steps performed by the agent during a reaching movement are the following: (i) the 0th order imposes a trajectory to the 1st order and generates a sensory prediction; (ii) the 0th order infers the consequences of its predictions, hence it is now biased toward both the intentional state and the observation; (iii) a proprioceptive prediction is generated from this new biased position, eventually driving action. This approach can be seen as a generalization of [63] where, in the context of oculomotor behavior, the target and the center of gaze were encoded as hidden states, each with their own dynamics and attracted by a hidden location. Although limited compared to non-linear dynamics functions (e.g., obstacle avoidance can be realized via repulsive potentials [80]), with the specific form defined above – along with the hidden states factorization – there is a high flexibility for complex interactions. Further, interpreting the hidden causes as a gain still makes sense from an active inference perspective because what is represented at a higher level is the intention to move at the target, while the target location is inferred at a lower level.
@@
Taken alone, considering a hidden cause as an attractor gain may not seem so helpful. However, as depicted in Figure 5(a), we can combine $M$ intentions in the following way:
$\begin{split}\bm{\eta}_{x}^{\prime}&=\bm{f}(\bm{x},\bm{v})=\sum_{m}^{M}\bm{f}_{m}(\bm{x},v_{m})=\sum_{m}^{M}v_{m}(\bm{i}_{m}(\bm{x})-\bm{x})\end{split}$ (15)
@@
In short, the agent entertains $M$ distinct dynamic hypotheses of how the world may evolve. Trajectories $\bm{f}_{m}(\bm{x},v_{m})$ are separately computed from each intention $\bm{i}_{m}$ and with their respective gains; then, the final trajectory $\bm{\eta}_{x}^{\prime}$ is found by combining all of them. The reason why we used the prior notation for the trajectory predictions will be clear in Chapter 4. Since each attractor $\bm{e}_{i,m}$ is proportional to its hidden cause $v_{m}$, the latter lends itself to a parallelism with the policies of discrete models, as will be explained later: if $v_{m}$ is set to $1$ and all the others to 00, the hidden states will be subject only to intention $m$; conversely, if multiple hidden causes are active, the hidden states will be pulled toward a combination of the corresponding intentions. This means that the hidden causes act both as attractor gains – expressing the absolute strength by which the belief is steered toward the desired direction – and as intention modulators – defining the relative strength between each trajectory. As a result, we have an additional modulation that combines with the dynamics precision $\bm{\Pi}_{x}$; their interactions will be explained in Chapter 4. The resulting behavior is similar to the Lotka-Volterra dynamics used in [81]. In the latter, the agent revisits in sequence points in attractor space linked to specific locations, and the defined dynamics ensure that only one attractor is active at any time. This permits modeling itinerant movements such as handwriting or walking.